OpenStudy (nathanjhw):

For the function f(x)=x^2-4x+5, x is greater than or equal to 2, which is equal to d/dx (f^(-1)(x))

OpenStudy (nathanjhw):

1/(2y-4) where x and y are related by the equation (satisfy the equation) x=y^2-4y+5, x is greater than or equal to 1 2y-4, where x and y are related by the equation y=x^2-4x+5, x is greater than or equal to 2 1/(2x-4), for x is greater than or equal to 1 1/(2x-4), for x is greater than or equal to 2 1/(2y-4), where x and y are related by the equation y=x^2 -4x +5, x is greater than or equal to 2

OpenStudy (freckles):

ok you need to know how to solve a quadratic equation I prefer completing the square when it comes to inverse of a quad \[y=x^2-4x+5 \\ y=x^2-4x+4+1 \\ y-1=x^2-4x+4 \] solve for x first complete the square

OpenStudy (nathanjhw):

x = sqrt(y-1) +2

OpenStudy (freckles):

right \[y-1=(x-2)^2 \\ x>2 \text{ \implies } x-2>0 \\ \text{ so } \sqrt{y-1}=x-2 \\ 2+\sqrt{y-1}=x \\ f^{-1}(x)=\sqrt{x-1}+2 \]

OpenStudy (freckles):

now differentiate

OpenStudy (nathanjhw):

(d/dx) sqrt(x-1) +2

OpenStudy (nathanjhw):

correct?

OpenStudy (freckles):

\[\frac{d}{dx}(f^{-1}(x))=\frac{d}{dx}(\sqrt{x-1}+2)\]

OpenStudy (nathanjhw):

1/f = 1/2(sqrt(x-1))

OpenStudy (freckles):

you mean the derivative right?

OpenStudy (freckles):

I don't know where 1/f comes from

OpenStudy (nathanjhw):

Yes

OpenStudy (astrophysics):

Inverse is not 1/f, a lot of people confuse that

OpenStudy (astrophysics):

\[f^{-1}(x) \neq \frac{ 1 }{ f(x) }\]

OpenStudy (nathanjhw):

Ah I see.

OpenStudy (freckles):

\[x=2+\sqrt{y-1} \\ \text{ we are already have } x \ge 2 \text{ and we also have that we need } y \ge 1 \] \[\frac{d}{dx} (f^{-1}(x))=\frac{1}{2 \sqrt{x-1}} \text{ where } x>1\]

OpenStudy (astrophysics):

Great explanations @freckles

OpenStudy (freckles):

thanks kindly

OpenStudy (nathanjhw):

Is that the final step because that is not one of my possible answer choices.

OpenStudy (freckles):

we switched x and y right? think we had y=x^2-4x+5 but here x is what? x=y^2-4y+5

OpenStudy (freckles):

if you replace x with that what do you have

OpenStudy (freckles):

\[\frac{1}{2 \sqrt{x-1}}=\frac{1}{2 \sqrt{y^2-4y+5-1}}=?\]

OpenStudy (freckles):

remember since we switched x and y we have x>1 and y>2 so y-2>0

OpenStudy (freckles):

you will need that y-2>0 to simplify

OpenStudy (nathanjhw):

So if we have x =y^2-4y+5 where x is greater than or equal to 1. We get the answer:1/(2y-4) where x and y are related by the equation (satisfy the equation) x=y^2-4y+5, x is greater than or equal to 1

OpenStudy (freckles):

close \[\frac{1}{2 \sqrt{x-1}}=\frac{1}{2 \sqrt{y^2-4y+4+1-1}}=\frac{1}{2 \sqrt{(y-2)^2}} =\frac{1}{2 (y-2)} \text{ since } y>2 \] and x>1 not equal to 1 since you can't plug 1 into that one expression thought your question asked for something in terms of x not y

OpenStudy (freckles):

so again you have to change the variables around again :p

OpenStudy (freckles):

just interchange them replace y with x

OpenStudy (nathanjhw):

So which answer choice would you say is right?

OpenStudy (freckles):

don't forget to change the one in the inequality too

OpenStudy (freckles):

the 4th option is what I'm implying

OpenStudy (nathanjhw):

Wow that question was confusing. I'm going to spend some time looking over this and see if I can understand it better. Thanks for all your help!

OpenStudy (freckles):

We did a lot of back and forth with the variables

OpenStudy (freckles):

there was something else we could have looked at

OpenStudy (freckles):

\[(f^{-1})'(x)=\frac{1}{f'(f^{-1}(x))}\] do you know this formula?

OpenStudy (freckles):

\[x^2-4x+5=y \\ \ (x-2)^2+1=y \\ (x-2)^2=y-1 \\ x-2=\sqrt{y-1} \text{ since } x \ge 2 \\ \text{ this also says } y \ge 1 \\ x=\sqrt{y-1}+2 \\ \text{ interchange } \\ y=\sqrt{x-1}+2 , y \ge 2 , x \ge 1 \\ f^{-1}(x)=\sqrt{x-1}+2 \\ \text{ now } f(x)=x^2-4x+5 \\ \text{ so } f'(x)=2x-4 \\ f'(f^{-1}(x))=2(\sqrt{x-1}+2)-4 =2 \sqrt{x-1} \\ \text{ so } (f^{-1})'(x)=\frac{1}{f'(f^{-1}(x))}=\frac{1}{2 \sqrt{x-1}}, x>1 , y>2 \] you know what I'm kinda confused about your choices again

OpenStudy (freckles):

@rational what do you think of these choices?

OpenStudy (freckles):

or @Astrophysics

OpenStudy (freckles):

maybe you are right the first time i thought wasn't it right because we didn't have a function of x

OpenStudy (rational):

I always end up with deriving the derivative of inverse \[f^{-1}(f(x)) = x \\~\\ [f^{-1}(f(x))]' f'(x)= 1\\~\\ [f^{-1}(f(x))]' = \dfrac{1}{f'(x)} \]

OpenStudy (rational):

but that works nicely for a specific value of x i think what you did is exactly is what is asked by the question i think as they are asking for x >= 2

OpenStudy (freckles):

@NathanJHW the more I think of the 1st one is right I think I'm going to say is right because it it telling you where x=y^2-4y+5 like it is telling you x-1=(y-2)^2 or x=(y-2)^2+1

OpenStudy (freckles):

still I think these choices are extremely odd :p

OpenStudy (freckles):

lol I made it harder that it should be

OpenStudy (rational):

x = y^2-4y+5 dx/dy = 2y - 4 dy/dx = 1/(2y-4) we still need to plugin the value of inverse function y here

OpenStudy (freckles):

like they didn't actually solve for the inverse though is the crazy thing

OpenStudy (rational):

i dont think so, solving the inverse function is still required

OpenStudy (freckles):

I did more work than needed

OpenStudy (freckles):

1/(2y-4) where x and y are related by the equation (satisfy the equation) x=y^2-4y+5, x is greater than or equal to 1 is the first choice this choice did not go as far as solving for the y given x=y^2-4y+5

OpenStudy (rational):

oh ok they gave options gotcha :)

OpenStudy (freckles):

thanks @rational for your workings of the problem :)